Hello, can you please help me solve this problem? I've asked mathnerds and read purple math and no one can help me solve it. I entered the problem here but it was returned stating No functions. The question is:
If x does not equal 0, then (2^4x)(8^2x) =
Thanks for your help, Susan

Found 2 solutions by stanbon, timmy1729:
Answer by stanbon(75887)

If x does not equal 0, then (2^4x)(8^2x)
2^(4x) * (2^3)^2x
= 2^(4x) *2^(6x)
=2^(10x)
= (2^10)(2^x)
= 1024*2^x
==============
Cheers,
Stan H.


Answer by timmy1729(23)

If x does not equal 0, then (2^4x)(8^2x) =
(2^4x)(2^3)^2x =
(2^4x)(2^6x) =
2^(4x+6x) =
2^(10x) =
(2^10)^x = 1024^x